# RD2Bench: Toward Data-Centric Automatic R&D

### v1

@@ 3.1 Experimental Settings
As we have numeric input and output in R&D tasks, we set numeric equability with 1e-6 as tolerance for the evaluation of the implementation of methods. We set the base models as GPT-4-turbo (OpenAI, 2023a), GPT-4-32k (OpenAI, 2023a), GPT-35-turbo-16k (OpenAI, 2023a) and Llama2 (Touvron et al., 2023) for the experiments. All the methods mentioned above, and their corresponding results are executed with Azure OpenAI API. There is no external data, resources, human feedback, or internet access involved in the experiments. We perform 20 independent attempts for each model and calculate the average and maximum value of each metric.
@@ 3.2 Results of Formula Implementation
In this section, we compare the performance of different models in the model architecture implementation task. We use the proposed metrics to evaluate the performance of the models. The results are shown in Table 1, Table 2, and Table 3. We observe that the GPT-4-turbo and GPT-4-32k achieve better performance than GPT-35-turbo-16k and LLaMa-2-70b in the model architecture implementation task. Overall experimental results indicate ample room for further research and the difficulty of the task and the challenges in automating R&D tasks. Specifically, we obtain the following four major findings revealed by the experimental results.
Category Difficulty Method Name avg. exe. avg. form. avg. corr. max. corr. max. acc. avg. acc. Data I simple PEG 90.00% 0.00% None None None None medium Turnover_STD_1M 80.00% 60.00% 0.314 0.839 0.001 0.000 hard turnover_correlation_with_price 20.00% 10.00% 0.371 0.936 0.000 0.000 discovery Liquidity_Factor 90.00% 55.00% None None None None Data II simple One_Month_Volatility 75.00% 45.00% 0.770 1.000 0.000 0.000 medium Vol20 90.00% 15.00% 0.566 1.000 0.000 0.0 hard Alpha#70 15.00% 0.00% None None None None Data III simple DailyRDvar 20.00% 0.00% None None None 0.005 medium AdjRDvar 65.00% 0.00% None None None 0.003 hard AdjRDskew 40.00% 5.00% NaN None None 0.000 discovery minute_pv_corr 35.00% 5.00% 0.915 0.947 0.744 0.248
Table 1: The performance of GPT-4-turbo on formula implementation task.
Category Difficulty Method Name avg. exe. avg. form. avg. corr. max. corr. max. acc. avg. acc. Data I simple PEG 75.00% 0.00% None None None None medium Turnover_STD_1M 75.00% 75.00% 0.194 0.243 0.000 0.000 hard turnover_correlation_with_price 25.00% 15.00% 0.123 0.193 0.000 0.000 discovery Liquidity_Factor 85.00% 50.00% None None None None Data II simple One_Month_Volatility 60.00% 40.00% 0.474 1.000 0.000 0.000 medium Vol20 20.00% 20.00% 0.527 1.000 0.000 0.000 hard Alpha#70 20.00% 0.00% None None None None Data III simple DailyRDvar 35.00% 5.00% None None 0.008 0.001 medium AdjRDvar 35.00% 0.00% None None 0.001 0.000 hard AdjRDskew 40.00% 5.00% NaN None None 0.000 discovery minute_pv_corr 20.00% 0.00% 0.434 0.858 0.002 0.001
Table 2: The performance of GPT-4-32k on formula implementation task.
@@
LLM agents hold promising potential to tackle D-CARD. We can observe from Table 1 and Table 2 that GPT-4 possesses the ability to tackle some simple D-CARD cases without adopting any additional techniques. Specifically, GPT-4 achieves a high (more than 0.8) maximum correlation coefficient with the ground-truth results in implementing both simple and medium formulations. Especially, both GPT-4-turbo and GPT-4-32k achieve the maximum correlation value in implementing simple One_Month_Volatility and Vol20, respectively. However, GPT-4 fails to precisely match the exact ground-truth values due to some minor mistakes, such as missing the domain common knowledge (e.g., using percent change rather than using difference when calculating the growth), mismatching the output format, and unnecessarily introducing additional computational operations.
@@
Precisely understanding and selecting data requires more detailed data information in D-CARD. As shown in Table 1 and Table 2, we observe a special situation where GPT-4 significantly fails to implement a simple formulation while succeeds in implementing the harder ones. After analyzing its generated code, we find that GPT-4 confuses the different semantic meanings of data features due to their close natural language descriptions, which renders the subsequent calculation ineffective. For example, GPT-4 confuses the two terms named “volume” and “volatility” and always opts to use “volume” data when “volatility” is required. If we manually improve our initial prompt by adding a more detailed description, GPT-4 succeeds in understanding the semantic difference and obtains over 99% performance in the accuracy of values.
@@
The ability to query domain-specific knowledge is a basic requirement of D-CARD methods. As we mentioned in the first finding, missing domain common knowledge impedes GPT-4 from calculating the precisely matched final results. Additionally, we find that the implementation of some operations in a formulation also requires domain-specific knowledge. For example, in the financial domain, it’s clear enough for the financial practitioners to implement the operation named “IndNeutralize(x,g)” by merely giving the description “x cross-sectionally neutralized against groups g”. However, in the code generated by GPT-4, it defines a function named “IndNeutralize(series, industry)” and leaves its content blank by merely adding a notation “Please replace this with your actual function definition”.
@@
The more complex the method is, the more unstable the model performance is. As shown in the columns of Table 1 and Table 2 named “avg. exe.”, “avg. form.”, and “avg. corr.”, respectively, we can observe that the performance variance of GPT-4 is significantly higher with the complexity of formulations getting higher. In 20 times of execution, GPT-4 generates the successfully executed code in 18 times when implementing the medium Vol20 while only three times in implementing hard Alpha#70.
Category Difficulty Method Name GPT-35-turbo-16k LLaMa-2-70b avg. exe. avg. form. avg. exe. avg. form. Data I simple PEG 0.00% 0.00% 0.00% 0.00% medium Turnover_STD_1M 0.00% 0.00% 0.00% 0.00% hard turnover_correlation_with_price 0.00% 0.00% 0.00% 0.00% discovery Liquidity_Factor 0.00% 0.00% 0.00% 0.00% Data II simple One_Month_Volatility 0.00% 0.00% 0.00% 0.00% medium Vol20 0.00% 0.00% 0.00% 0.00% hard Alpha#70 0.00% 0.00% 0.00% 0.00% Data III simple DailyRDvar 0.00% 0.00% 0.00% 0.00% medium AdjRDvar 0.00% 0.00% 0.00% 0.00% hard AdjRDskew 5.00% 0.00% 0.00% 0.00% discovery minute_pv_corr 5.00% 0.00% 0.00% 0.00%
Table 3: The performance of GPT-35-turbo-16k and LLaMa-2-70b on formula implementation task.
@@
As shown in Table  3, the performance of GPT-35 and LLaMa2 is poor, even failing in execution codes. However, GPT-4 models shown in Table  1 and  2 have a much better performance. This indicates that the performance of the model in the data-centric R&D task is highly related to the model’s pre-training and capacity. Therefore, we posit that continually training and improving the foundation model is a promising direction for future research in the field of data-centric R&D tasks.
@@ 3.3 Results of Model Architecture Implementation
In this section, we compare the performance of different LLMs in the model architecture implementation task and summarize the results in Table  4 and  5. As shown in the table, we can see the GPT-4-turbo, GPT-35-turbo-16k, and GPT-4-32K have similar running success rates, but differ variously in tvc. and tsc.. The LLaMa-2-70b has the lowest running success rate and other metrics. Notice that even though a significant gap still exists between GPT-35, LLaMa-2, and GPT-4, it is much smaller than the gap in the formula implementation task. The overall running success rates are also higher than formula implementation task. We can conclude that we can have similar observations in the model architecture implementation task as in the formula implementation task.
@@
Table 4: The performance of GPT-4-turbo and GPT-4-32k on model architecture implementation task.
@@
Model Name Difficulty GPT-35-turbo-16k LLaMa-2-70b avg. exe. avg. tsc. avg. tvc. max. tsc. max. tvc. avg. exe. avg. tsc. avg. tvc. max. tsc. max. tvc. PMLP Easy 100.00% 0.75 0.75 1.00 1.00 60.00% 0.45 0.55 1.00 1.00 LinkX Easy 100.00% 0.60 0.34 1.00 1.00 30.00% 0.20 0.15 1.00 1.00 VisNet Hard 5.00% 0.03 0.00 0.16 0.40 0.00% 0.00 0.00 0.00 0.00 AntiSymmetric Medium 45.00% 0.16 0.21 0.61 0.22 0.00% 0.00 0.00 0.00 0.00 GPSConv Medium 45.00% 0.24 0.19 0.45 0.42 0.00% 0.00 0.00 0.00 0.00 DirGNNConv Medium 65.00% 0.56 0.29 0.71 0.42 0.00% 0.00 0.00 0.00 0.00
@@
Table 5: The performance of GPT-35-turbo-16k and LLaMa-2-70b on model architecture implementation task.
3.4 Limitations
We also strongly want to include as many as possible of current state-of-the-art evolving tricks and techniques in the benchmark in further works, which will lead to a more comprehensive evaluation and reveal more signals of the performance of the models. We have primarily tried to apply tricks like “self-refine/correct/reflection” and “automatic curriculum” to improve the performance of LLM agents in this benchmark, and we find that the improvement is marginal in the performance of the models. For more details and a more comprehensive evaluation, we will include them in not-far-future works. We believe that the benchmark will be a valuable tool for the community to evaluate the performance of the models in the data-centric R&D tasks and to develop new models and techniques to address the challenges and opportunities in the domain.
4 Relate Work
4.1 LLM as Autonomous Agent
In the past few years, LLM has made great achievements in both academia and industry (OpenAI, 2023a; Touvron et al., 2023), and has achieved results that surpass the previous level in a number of classic tasks (Zhao et al., 2023). Research has shown that with the growth of data volume and model size (Zoph et al., 2022), LLM has emerged with stronger reasoning and other capabilities (Ouyang et al., 2022). These capabilities enable LLM to exhibit certain agent-like behaviors in some tasks such as using or creating tools (Qin et al., 2023b; Qian et al., 2023), planning (Yao et al., 2023; Brown et al., 2020), and memory. Therefore, more and more researchers have expressed their expectations for its human-like and overall capabilities, and have made preliminary explorations of it as an independent agent (Wang et al., 2023a; Shinn et al., 2023). Multi-agent collaboration (Wu et al., 2023; Li et al., 2023) is also introduced to LLM for better accuracy and generalizability. Moreover, for reducing human efforts and automatically exploring, previous work focuses on autonomous LLM agents for general purpose are purposed (Yang et al., 2023b; Shen et al., 2023). Positive views further believe that the realization of AGI may come from the evolution of autonomous LLM and some inspiring examples have been released (Penov et al., 2024).
However, most research still focuses on limited scenarios that are given with clear and fixed questions and backgrounds. A recent work (Yang et al., 2023d) has attempted to introduce LLM to the R&D field and formalize the R&D process as a sequence of tasks. However, there is no easy-to-use benchmark for the community and current R&D tasks may be too general and can’t reveal significant signals. In this work, we propose a benchmark for LLM in data-centric R&D tasks and provide a comprehensive evaluation.
4.2 Semi-Automatic R&D with Agents
Scientific research and development (R&D) is a time-consuming and important process. In the past, R&D has been mainly conducted by human researchers with countless failed experimental explorations and creative observation conclusions. Agents have been introduced to R&D to reduce human efforts and automatically explore. Recently, there have been attempts to partly automate R&D, including the automatic chemical synthesis planning  (Boiko et al., 2023), automatic molecular design  (Joshi & Kumar, 2021; Schneider, 2017; Boiko et al., 2023), automatic theorem proving  (Wang et al., 2023b; Yang et al., 2023c). However, these attempts mainly focus on automatic searching for possible solutions and optimizations with symbolic representation  (Lu et al., 2023) and heuristic techniques  (Whalen, 2016), but less addressing long-horizon planning, implementation, and reasoning for the next step idea exploration. Moreover, the data-centric R&D tasks currently have not been explored in the community, and no benchmark has been proposed for the community. Previous works have applied LLM to real-world R&D tasks such as debugging issues  (Tian et al., 2024; Jimenez et al., 2023a) or only focus on data-centric but not real-world R&D tasks  (Liu et al., 2023a). In this work, we propose a benchmark for LLM in data-centric R&D tasks and evaluate the performance of LLMs.
@@ 5 Conclusion
In this paper, we serve as the first effort to tackle the real-world data-centric automatic R&D scenario in the hope of significantly improving the research efficiency of scientists and thus contributing to the revolution of human productivity. Specifically, we first propose RD2Bench that benchmarks all the operations in D-CARD as a whole to navigate future work toward the ultimate goal of automating data-centric R&D directly. RD2Bench focuses on evaluating the interaction and synergistic effects of various model capabilities and aiding in selecting the well-performing trustworthy models. Based on RD2Bench, we find that although the most SOTA GPT-4 shows its promising potency in tackling D-CARD, there remains ample room for future work.
Appendix A Appendix
A.1 Formula implementation task metrics calculation details
@@
Running success is defined as successful execution. Any error occurs in python interpreter during the execution that stop the execution is considered as a failure. We calculate the ratio of the number of successful execution times to the total number of attempts, denoted as avg. exe.
@@
Format success is defined as successful format matching, which is the final output dataframe format is (datetime, factor_name). We calculate the ratio of the number of the matched result formats to the total number of attempts, denoted as avg. form.
@@
Please note that we set the tolerance $t$ for the value accuracy as 1e-6 in this paper, which means two value is considered as equal if the absolute difference is less than 1e-6.
A.2 Data collection details
As mentioned in the previous section, we collect papers including  (Gravina et al., 2023; Rossi et al., 2023; Rampášek et al., 2022; Lim et al., 2021; Yang et al., 2023a; Wang et al., 2024) and corresponding codes using pyg (Fey & Lenssen, 2019), which are listed in the following table.

### v2

@@ 4.1 Experimental Settings
As we have numeric input and output in R&D tasks, we set numeric equability with 1e-6 as tolerance for the evaluation of the implementation of methods. We set the base models as GPT-4-turbo [22], GPT-4-32k [22], GPT-35-turbo-16k [22] and Llama2 [42] for the experiments. All the methods mentioned above, and their corresponding results are executed with Azure OpenAI API. There is no external data, resources, human feedback, or internet access involved in the experiments. We perform 20 independent attempts for each model and calculate the average and maximum value of each metric. As most of our input data is encoded in the form of document files, we first use parsing tools to extract text content from files. Azure document intelligence API (4.0) is used for parsing reports and academic papers in PDF format.
4.2 Results of Method Extraction
We evaluate the information extraction ability of models. As shown in Table 1, we observe that GPT-4-turbo, GPT-4-32k, LLaMa3-70b, and LLaMa2-70b achieve competitive performance, which makes them possible to perform information extraction automatically. The performance of the four foundation models is stronger than of Phi3-4k, indicating more future endeavors to improve the extraction ability of Phi3-4k.
Metrics Precion Recall F1
GPT-4-turbo 0.818 0.800 0.809
GPT-4-32k 0.818 0.818 0.818
LLaMa3-70b 0.909 0.833 0.869
LLaMa2-70b 0.818 0.900 0.857
Phi3-4k 0.636 0.750 0.688
Table 1: Results of method extraction.
@@ 4.3 Results of Formula Implementation
In this section, we compare the performance of different models in the model architecture implementation task. We use the proposed metrics to evaluate the performance of the models. The results are shown in Table 2 and Table 3. We observe that the GPT-4-turbo achieves better performance than GPT-35-turbo and Phi3-128k in the model architecture implementation task. Overall experimental results indicate ample room for further research on the difficulty of the task and the challenges in automating R&D tasks. Specifically, we obtain the following four major findings revealed by the experimental results.
Data Difficulty Formula Avg. Exec. Avg. Format Avg. Corr. Max. Corr.
Data I Easy PB_ROE 0.650 0.050 0.852 0.852
PB_ROE_2 0.600 0.200 0.875 1.000
PB_ROE_3 0.600 0.300 0.726 1.000
Medium ROE_movement 0.950 0.750 0.934 1.000
ROE_movement_10 0.900 0.800 0.803 1.000
ROE_movement_20 0.950 0.750 0.703 1.000
Hard PB_ROE_movement 0.600 0.450 0.516 0.897
PB_ROE_movement_10 0.650 0.300 0.327 0.896
PB_ROE_movement_20 0.550 0.500 0.244 0.896
Data II Easy mid_price 0.800 0.100 1.000 1.000
mid_price_2 0.850 0.000 NaN NaN
mid_price_3 0.850 0.000 NaN NaN
Medium liquidity_imbalance 0.500 0.050 1.000 1.000
liquidity_imbalance_2 0.900 0.150 0.694 1.000
liquidity_imbalance_3 0.450 0.100 1.000 1.000
Hard micro_price 0.850 0.000 NaN NaN
micro_price_2 0.600 0.000 NaN NaN
micro_price_3 0.600 0.100 1.000 1.000
Data III Easy alpha053 0.950 0.700 0.933 1.000
alpha053_15 0.950 0.650 0.872 1.000
alpha053_5 1.000 0.650 0.676 1.000
Medium alpha_pv_diff 1.000 0.600 0.513 1.000
alpha_pv_diff_15 0.950 0.750 0.258 1.000
alpha_pv_diff_20 1.000 0.750 0.441 1.000
Hard alpha_pv_diff_pct 0.950 0.700 0.375 1.000
alpha_pv_diff_pct_15 0.900 0.450 0.236 1.000
alpha_pv_diff_pct_20 1.000 0.350 0.358 1.000
Overall N/A Avg. Data I 0.717 0.456 0.665 0.949
Avg. Data II 0.711 0.056 0.522 0.556
Avg. Data III 0.967 0.622 0.518 1.000
Mean Value 0.798 0.378 0.568 0.835
Table 2: The performance of GPT-4-turbo in formula implementation.
@@
LLM agents hold promising potential to tackle D-CARD. We can observe from Table 2 and Table 3 that GPT-4 possesses the ability to tackle some simple D-CARD cases without adopting any additional techniques. Specifically, GPT-4 achieves a high maximum correlation coefficient with the ground-truth results in implementing both easy and medium formulations: GPT-4-turbo achieves the maximum correlation value in implementing easy and medium formulas. However, GPT-4 fails to precisely match the exact ground-truth values due to some minor mistakes, such as missing the domain common knowledge (e.g., using percent change rather than difference when calculating growth), mismatching the output format, and unnecessarily introducing additional computational operations.
@@
Precisely understanding and selecting data requires more detailed data information in D-CARD. As shown in Table 2, we observe a special situation where GPT-4 significantly fails to implement a simple formulation while succeeding in implementing the harder ones. After analyzing its generated code, we find that GPT-4 confuses the different semantic meanings of data features due to their close natural language descriptions, which renders the subsequent calculation ineffective. For example, GPT-4 confuses the two terms named “volume” and “volatility” and always opts to use “volume” data when “volatility” is required. If we manually improve our initial prompt by adding a more detailed description, GPT-4 succeeds in understanding the semantic difference and obtains over 99% performance in the accuracy of values.
@@
The ability to query domain-specific knowledge is a basic requirement of D-CARD methods. As we mentioned in the first finding, missing domain common knowledge impedes GPT-4 from calculating precisely matched final results. Additionally, we find that the implementation of some operations in a formulation also requires domain-specific knowledge. For example, in the financial domain, it’s clear enough for financial practitioners to implement the operation named “IndNeutralize(x,g)” by merely giving the description “x cross-sectionally neutralized against groups g”. However, in the code generated by GPT-4, it defines a function named “IndNeutralize(series, industry)” and leaves its content blank by merely adding a notation “Please replace this with your actual function definition”.
@@
The more complex the method is, the more unstable the model performance is. As shown in the columns of Table 2 named “avg. exec.”, “avg. form.”, and “avg. corr.”, respectively, we can observe that the performance variance of GPT-4 is significantly higher as the complexity of formulations increases. In 20 times of execution, GPT-4 generates the successfully executed code 18 times when implementing the medium mid_price while only three times in implementing hard alpha_pv.
avg. exec. avg. format avg. corr. max. corr.
GPT-4—turbo Data I 0.717 0.456 0.665 0.949
Data II 0.711 0.056 0.522 0.556
Data III 0.967 0.622 0.518 1.000
Mean Value 0.798 0.378 0.568 0.835
GPT-35-turbo Data I 0.556 0.100 0.323 0.453
Data II 0.567 0.000 0.000 0.000
Data III 0.767 0.389 0.431 0.696
Mean Value 0.630 0.163 0.251 0.383
Phi3-128k Data I 0.117 0.111 0.186 0.222
Data II 0.172 0.000 0.000 0.000
Data III 0.056 0.022 0.063 0.084
Mean Value 0.115 0.044 0.083 0.102
Table 3: Comparison of our benchmark for different models
@@
As shown in Table  3, the performance of GPT-35-turbo and Phi3-128k is poor, even failing in execution codes. However, GPT-4 models shown in Table 2 have a much better performance. This indicates that the performance of the model in the data-centric R&D task is highly related to the model’s pre-training and capacity. Therefore, we posit that continually training and improving the foundation model is a promising direction for future research in the field of data-centric R&D tasks.
@@ 4.4 Results of Model Architecture Implementation
In this section, we compare the performance of different LLMs in the model architecture implementation task and summarize the results in Table  4 and  5. As shown in the table, we can see the GPT-4-turbo, GPT-35-turbo-16k, and GPT-4-32K have similar running success rates but differ variously in tvc. and tsc.. The LLaMa-2-70b has the lowest running success rate and other metrics. Notice that even though a significant gap still exists between GPT-35, LLaMa-2, and GPT-4, it is much smaller than the gap in the formula implementation task. The overall running success rates are also higher than the formula implementation task. We can conclude that we can have similar observations in the model architecture implementation task as in the formula implementation task.
@@
Table 4: The performance of GPT-4-turbo and GPT-4-32k on model architecture implementation tasks.
@@
Model Name Difficulty GPT-35-turbo-16k LLaMa-2-70b LLaMa-3-70b avg. exe. avg. tsc. avg. tvc. max. tsc. max. tvc. avg. exe. avg. tsc. avg. tvc. max. tsc. max. tvc. avg. exe. avg. tsc. avg. tvc. max. tsc. max. tvc. PMLP Easy 100.00% 0.75 0.75 1.00 1.00 60.00% 0.45 0.55 1.00 1.00 100.00% 0.85 0.75 1.00 1.00 LinkX Easy 100.00% 0.60 0.34 1.00 1.00 30.00% 0.20 0.15 1.00 1.00 100.00% 0.80 0.65 1.00 1.00 VisNet Hard 5.00% 0.03 0.00 0.16 0.40 0.00% 0.00 0.00 0.00 0.00 40% 0.27 0.24 0.33 0.42 AntiSymmetric Medium 45.00% 0.16 0.21 0.61 0.22 0.00% 0.00 0.00 0.00 0.00 80% 0.62 0.70 0.71 0.88 GPSConv Medium 45.00% 0.24 0.19 0.45 0.42 0.00% 0.00 0.00 0.00 0.00 75% 0.51 0.59 0.65 1.00 DirGNNConv Medium 65.00% 0.56 0.29 0.71 0.42 0.00% 0.00 0.00 0.00 0.00 90% 0.72 0.68 0.84 0.91
@@
Table 5: The performance of GPT-35-turbo-16k, LLaMa-2-70b and LLaMa-3-70b on model architecture implementation tasks.
5 Limitation
The $\text{RD}^{2}$ench framework, while innovative, only evaluates the most representative base LLM, such as GPT4, LLama3, LLama2, GPT35, without further evaluate more open source models. Meanwhile, this paper only include most representative R&D domains and problems and focus on data driven scenarios, which can be extended more in the future to show its generalizability. For more details and a more comprehensive evaluation, we will include them in not-too-distant future works. We believe that the benchmark will be a valuable tool for the community to evaluate the performance of the models in the data-centric R&D tasks and to develop new models and techniques to address the challenges and opportunities in the domain.
@@ 6 Conclusion
In this paper, we serve as the first effort to tackle the real-world data-centric automatic R&D scenario in the hope of significantly improving the research efficiency of scientists and thus contributing to the revolution of human productivity. Specifically, we first propose $\text{RD}^{2}$ench that benchmarks all the operations in D-CARD as a whole to navigate future work toward the ultimate goal of automating data-centric R&D directly. $\text{RD}^{2}$ench focuses on evaluating the interaction and synergistic effects of various model capabilities and aiding in selecting the well-performing trustworthy models. Based on $\text{RD}^{2}$ench, we find that although the most SOTA GPT-4 shows its promising potency in tackling D-CARD, there remains ample room for future work.
Appendix A Formula implementation task metrics calculation details
@@
Running success is defined as successful execution. Any error that occurs in the Python interpreter during the execution that stops the execution is considered a failure. We calculate the ratio of the number of successful execution times to the total number of attempts, denoted as avg. exe.
@@
Format success is defined as successful format matching, which means the final output dataframe format is (datetime, factor_name). We calculate the ratio of the number of matched result formats to the total number of attempts, denoted as avg. form.
@@
Please note that we set the tolerance $t$ for the value accuracy as 1e-6 in this paper, which means two values are considered as equal if the absolute difference is less than 1e-6.
Appendix B Data collection details
As mentioned in the previous section, we collected papers including  [10, 34, 32, 15, 50, 47] and corresponding codes using pyg [8], which are listed in the following table.